OpenStudy (anonymous):

A small rocket is launched with an acceleration of 10 m/s^2. the velocity, v, of the rocket after it has traveled s meters is given by v^2=20s. Find the velocity of this rocket after it has traveled 500m

OpenStudy (whpalmer4):

\[v^2=20s\]Find value of \(v\) when \(s = 500\text{ m}\) \[v^2=20*500\]Solve for \(v\)

OpenStudy (anonymous):

THANK YOU

OpenStudy (whpalmer4):

You're welcome! What did you get for your answer?

OpenStudy (anonymous):

V^2=10,000

OpenStudy (whpalmer4):

Okay, so far so good. But what is the value of \(v\)?

OpenStudy (whpalmer4):

We know what \(v*v\) is, but we want to know \(v\).